An investor bought a one-acre lot on the outskirts of a city for $12,700cash. Each year she paid $175 of property taxes . At the end of 7 years, she lost the plot for a net value of $25,000. What rate of return did she receive on her investment?
Answer #1

FV = PV(1+r)^n

r = Rate of ret

n = holding period

FV = PV(1+r)^n

$ 25,000 = $ 12,875 (1+r)^7

(1+r)^7 = 25000 / 12875

= 1.9418

1+r = (1.9418)^(1/7)

= 1.0994

r = 1.0994 - 1

= 0.0994 i.e 9.94%